OpenStudy (darkbluechocobo):

The following lines are parallel to each other. 2x − 5y = -3 and 5y + 2x = 20 Is this false?

OpenStudy (kkutie7):

solve each for y first.

OpenStudy (darkbluechocobo):

5y= 2x+3 5y=-2x+20 y=2x+3/5 and y=-2x+20/5

OpenStudy (kkutie7):

look at slope in front of the x. if they are the same they are parallel.